OpenStudy (anonymous):

estimate the derivative of f(x)=x^x at x=2

OpenStudy (mathteacher1729):

Estimate using which method?

OpenStudy (anonymous):

if it really means "estimate" whip out your calculator and try \[\frac{(2.001)^{2.001}-2^2}{.001}\]

OpenStudy (anonymous):

if you want to find the "actual value" you will either have to use logarithmic differentiation or else write \[f(x)=x^x=e^{x\ln(x)}\] and take the derivative using the chain rule and the product rule.

OpenStudy (anonymous):

for the example i did above i got 6.779 rounded. if you want a better estimate pick a decimal even closer to 2. for example try \[\frac{(2.000001)^{2.000001}-4}{.000001}\]

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

yw